OpenStudy (anonymous):

A recipe that makes 4 servings calls for 3/4 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving? A. 5/7 B. 17/48 C. 1 5/12 D. 5 2/3

OpenStudy (breanna77):

B

OpenStudy (anonymous):

how did you get the answer @breanna77

OpenStudy (breanna77):

3/4 + 2/3 = 9/12 + 8/12 = 17/12 (17/12) / 4 = (17/12) * (1/4) = 17/48 cup per serving

OpenStudy (anonymous):

ok thankyou @breanna77